OpenStudy (anonymous):

If a hospital patient is given 50 milligrams of medicine which leaves the bloodstream at 10% per hour, how many milligrams of medicine will remain in the system after 6 hours? Use the function A(t) = Iert. 12.39 mg 27.44 mg 45.24 mg 91.11 mg when i plug it in i get 50e^.6 and i find the answer is 91.11 but on other sites it's 27.44

OpenStudy (mathmath333):

the formula is \(A(t)=le^{rt}\\ =\) A=50,r=10,t=6. \(\Large 50=l\times e^{10\times 6}\\ \Large l=\dfrac{50}{e^{ 0.6}}\\ \)

OpenStudy (mathmath333):

wolfram does gives 27.44 , their is typo in ur calculation

OpenStudy (anonymous):

oh ok thank you so much , writing the formula down like you did and going through the problem really helped me to get it , thanks a lot

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

yw